Chapter 09 Homework Saved Required information Part 1 of 3 [The following information applies to the questions displayed beloChapter 09 Homeworki Saved Required information Part 2 of 3 [The following information applies to the questions displayed belEmployee Paid Таxes General Journal Assuming situation a, prepare the employers September 30 journal entries to record salarChapter 09 Homework Saved 10 ! Required information [The following information applies to the questions displayed below] PartEmployer Payroll Taxes General Journal Assuming situation a, prepare the employers September 30 journal entries to record th
Answer #1
a)
Tax September Earnings Subject to Tax Tax Rate Tax Amount
FICA-Social Security 2200 6.20% 136.40
FICA-Medicare 2200 1.45% 31.90
FUTA 1900 3.10% 58.90
SUTA 1900 2.90% 55.10
b)
Tax September Earnings Subject to Tax Tax Rate Tax Amount
FICA-Social Security 2700 6.20% 167.40
FICA-Medicare 2700 1.45% 39.15
FUTA 2700 3.10% 83.70
SUTA 2700 2.90% 78.30
c)
Tax September Earnings Subject to Tax Tax Rate Tax Amount
FICA-Social Security 6300 6.20% 390.60
FICA-Medicare 8600 1.45% 124.70
FUTA 0 3.10% 0.00
SUTA 0 2.90% 0.00
Note: Earnings taxable for FUTA and SUTA in (a) = 7000-5100 = $1900
Note: Earnings taxable for Social Security in (c) = 127200-120900 = $6300
2
Taxes to be Withheld From Gross Pay (Employee-Paid Taxes)
September Earnings Subject to Tax Tax Rate Tax Amount
Federal income tax 80.00
FICA—Social Security 2200 6.20% 136.40
FICA—Medicare 2200 1.45% 31.90
Total taxes withheld 248.30
Date General Journal Debit Credit
Sep 30 Salaries expense 2200.00
FICA—Social Security taxes payable 136.40
FICA—Medicare taxes payable 31.90
Employee federal income taxes payable 80.00
Salaries payable 1951.70
3
Employer Payroll taxes September earnings subject to tax Tax Rate Tax Amount
FICA-Social Security 2200 6.20% 136.40
FICA-Medicare 2200 1.45% 31.90
FUTA 1900 3.10% 58.90
SUTA 1900 2.90% 55.10
Total Payroll Taxes Expense 282.30
Date General Journal Debit Credit
Sep. 30 Payroll taxes expense 282.30
FICA—Social security taxes payable 136.40
FICA—Medicare taxes payable 31.90
Federal unemployment taxes payable 58.90
State unemployment taxes payable 55.10